OpenStudy (anonymous):

Look at the table of values below. x y 1 -1 2 -3 3 -5 4 -7 Which equation is represented by the table?

OpenStudy (anonymous):

A. y = 1 - 2x B. y = -x - 1 C. y = x - 2 D. y = 2x - 1

jimthompson5910 (jim_thompson5910):

if you plug x = 1 into the equation for choice A, what do you get? if you plug x = 2 into the equation for choice A, what do you get? if you plug x = 3 into the equation for choice A, what do you get? if you plug x = 4 into the equation for choice A, what do you get? If you get y values that match up with the table, then you have found your answer. If not, then move onto the next equation and repeat the last 4 steps.

OpenStudy (anonymous):

@jim_thompson5910 im just looking for the answer, not interested i solving right now

OpenStudy (anonymous):

in